Lisas tuna salad recipe calls for 4 tablespoons of mayonnaise for every 3lb of tuna fish. How many tablespoons of mayo does she need to make 10lb of tuna salad?
Answer by rfer(16322)

4/3=x/10
3x=40
x=40/3
x=13 1/3